[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Nov 09, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Mon Nov 14 17:43:22 UTC 2016


Here is today's summary of NHC computer operations:
--- Craig Mattocks

----------

TAFB:
1. The (center) security camera is not displaying the northwest single-door  
area on the WFO and TAFB monitors. This is probably due to a loose monitor  
cable because the security guards are viewing the live video feed without  
problems. An NCO Help Desk ticket will be filed.

2. The light fixture over the TAFB backup desk is flickering, even though  
the light bulb has been replaced recently. Maintenance personnel will be  
contacted.
HSU:

Other items:
1. NESDIS STAR's ftp2 server is on its last legs. A new ftp3 server has  
been set up and users are urged to begin pulling data from it as a test  
ASAP. NCO's Dataflow group has been advised of the situration.

NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161114/b5db546d/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list